# Continuously Learning New Words in Automatic Speech Recognition

### v2

@@ 4.1 Limitations
The extraction of words from the slides is non-trivial. We took all words from the slides and filtered by the words in the training data. Errors occur when the words on the slides are capitalized but in the transcript they should not be capitalized and when the extracted words contain spelling errors. A further problem is when morphological variances of known words are extracted, as their presence in memory increases the rate of false positives. Adding other morphological variances of the same word may help here. All these mentioned errors can propagate through the learning cycles. Moreover, the training with pseudo labels could lead to a degradation of the general performance. We did not observe that, however, one could restrict the training for the pseudo labels to tokens which are part of new words. Other false positives can occur, e.g., when first a word like ISCA is learned followed by a word ISA which is written similarly. In this case, a small supervised intervention could help to reduce the amount of false-positive utterances added to the adaptation data.
5 Conclusion
@@
We demonstrated an self-supervised continual learning approach for learning new words. This is done by iteratively extracting new words from slides of a given talk, detecting new words by a memory-enhanced ASR model and using collected data for adapting low-rank matrix weights added to each weight matrix of the model. With this approach, we can increase the performance of new words as they occur more often to more than 80 % recall while the general performance of the model is preserved.
6 Acknowledgements

### v3

@@ IV-A Limitations
The extraction of words from the slides is not trivial. We took all words from the slides and filtered by the words in the training data. Errors occur when the words on the slides are capitalized but in the transcript they should not be capitalized, and when the extracted words contain spelling errors. A further problem arises when the morphological variances of known words are extracted, as their presence in memory increases the rate of false positives. Adding other morphological variances of the same word may help here. All these mentioned errors can propagate through the learning cycles. Moreover, training with pseudo-labels could lead to a degradation of the general performance. We did not observe that, however, one could restrict the training for the pseudo-labels to tokens that are part of new words. Other false positives can occur, e.g., when first a word like ISCA is learned followed by a word ISA which is written similarly. In this case, a small supervised intervention could help reduce the number of false positive utterances added to the adaptation data.
V Conclusion
@@
We demonstrated a self-supervised continuous learning approach for learning new words. This is done by iteratively extracting new words from slides of a given talk, detecting new words by a memory-enhanced ASR model, and using collected data for adapting low-rank matrix weights added to each weight matrix of the model. With this approach, we can increase the performance of new words as they occur more often to more than 80 % recall while the general performance of the model is preserved.
VI Acknowledgements